OpenStudy (anonymous):

derivative of sqrt(2x^2 - 2x + 3) please

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\) so by the chain rule, the derivative of \[\sqrt{f(x)}\] is \[\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

in other words, take the derivative of \(2x^2-2x+3\) and put that in the numerator and the denominator will be \(2\sqrt{x^2-2x+3}\)

OpenStudy (anonymous):

sorry typo there the denominator will be \(2\sqrt{2x^2-2x+3}\)